Shanti is 5 times as old as her brother. Next year, she will be 3 times as old. What is Shanti's age now?
How do you solve this?

Answer by Alan3354(69443)

s = her age, b = brother's age
----
Now, s = 5b
Next year,
s+1 = 3(b+1) = 3b+3
----
s = 3b+2
s = 5b
Sub for s in the 1st equation
5b = 3b + 2
2b = 2
b = 1
------
s = 5